OpenStudy (anonymous):

I'm having trouble with figuring out whether a trinomial is square or not. How exactly do I do this?

OpenStudy (anonymous):

When you say square do you mean quadratic?

OpenStudy (anonymous):

I'm gonna assume that's what you mean. If the highest power of x in the equation is 2 then it's quadratic.

OpenStudy (anonymous):

Specifically; x squared - 14x = 49

OpenStudy (anonymous):

i think you may mean \[a^2+2ab+b^2=(a+b)^2\] yes?

OpenStudy (anonymous):

so if you see \[x^2-14x+49\] you think \[a=x\] \[b=-7\] because \[(-7)^2=49\] and \[2\times -7=-14\]

OpenStudy (anonymous):

and therefore \[x^2-14x+49=(x-7)^2\]